# Repeated Padding for Sequential Recommendation

### v2

@@ 3.8. Ablation Study
In this section, we conduct an ablation study to explore the different variants of RepPad. Based on the padding times $m$ in Table 1 and whether or not to add the delimiter 0, we compare the performance of the following variants: 1) RP-Fix: Performing repeated padding with a fixed number of times, we report the best results for times in range $\{1,2,3\}$. 2) RP-Max: Performing repeated padding until the remaining space is insufficient (i.e., maximum padding times). 3) RP(0, Max): When performing repeated padding each time, randomly select between 0 and the maximum number of times. 4) RP(1, Max): When performing repeated padding each time, randomly select between 1 and the maximum number of times. Figure 2 presents the results. We report the performance of each of the four variants with and without adding the delimiter 0.
@@
For performance between different variants, in most results, the performance of RP(0, Max) and RP(1, Max) can come with considerable improvements. RP-Fix and RP-Max bring limited performance improvement compared to RP(0, Max) and RP(1, Max). For PR-Fix, RepPad is executed with a fixed number of times, so it does not utilize the idle input space. For RP-Max, we observe that it not only does not bring improvement in some cases but also leads to a significant decrease in model performance. If every sequence performs RepPad to the maximum number of times, it may make the representation learned by the model unbalanced. Because the repeated padding time tends to be large for short sequences, multiple iterations of short behavioral sequences may interfere with the model’s ability for long-sequence and long-term preferences modeling, which in turn impairs recommendation performance. For RP(0, Max) and RP(1, Max), they avoid the imbalance problem caused by RP-Max while utilizing the free input space. Meanwhile, RP(1, Max) performs better than RP(0, Max), indicating that executing RepPad at least once is essential.
@@
For whether or not to add the delimiter 0, we observe that this factor is related to the type of recommendation model. For GRU4Rec, there are more examples of better performance when no delimiter is added. In contrast, for SASRec, there are more examples that models perform better when the delimiter is added. In other words, RNN-based models are less susceptible to the interference of ”predicting the tail with the head”. The self-attention module of Transformer-based models pays attention to the effect of each historical interaction on the following action, so the negative effect of ”predicting the tail with the head” can be minimized by adding the delimiter 0. We leave the theoretical explanation of this phenomenon to the future work.
@@
Table 8. The epoch time (seconds) of different methods. We give the increased time compared to the original model.
Dataset SASRec GRU4Rec
Method Epoch Time Method Epoch Time
Beauty Original 2.389 Original 0.960
w/ CMRSI 4.561 (+2.172) w/ CMRSI 2.547 (+1.587)
w/ CL4SRec 3.954 (+1.565) w/ CL4SRec 1.748 (+0.778)
w/ RP-Fix 2.483 (+0.094) w/ RP-Fix 1.049 (+0.089)
w/ RP(1, Max) 2.768 (+0.379) w/ RP(1, Max) 1.308 (+0.384)
Sports Original 3.133 Original 1.422
w/ CMRSI 6.400 (+3.267) w/ CMRSI 3.853 (+2.431)
w/ CL4SRec 5.479 (+2.346) w/ CL4SRec 2.620 (+1.198)
w/ RP-Fix 3.219 (+0.086) w/ RP-Fix 1.548 (+0.126)
w/ RP(1, Max) 3.740 (+0.607) w/ RP(1, Max) 1.998 (+0.576)
4. Why RepPad Works
4.1. Loss Convergence and Training Efficiency
In Figure 3, we illustrate the loss convergence of different methods during training. We also compare the epoch time consumption of different methods in Table 8. From the figures and tables, we can observe that RepPad’s losses converge slightly faster than the original model. For CMRSI, the heuristic data augmentation introduces more different training sequences along with more noise, resulting in more epochs for the model to converge. The contrastive learning task of CL4SRec also imposes additional computational overhead and convergence burden. Also, RepPad only adds a small amount of epoch elapsed time compared to the original model. While CMSRI and CL4SRec add extra training sequences and auxiliary tasks, resulting in a substantial increase in training elapsed time. Comparing RP-Fix and RP(1, Max), the former can only be performed repeated padding a fixed and smaller number of times. At the same time, the latter randomly chooses between one time and the maximum number of times, which slightly increases the elapsed time but improves the speed of loss convergence and model performance.
@@ 4.2. Gradient Stability
Research has shown that more stable or smooth gradients are more helpful for model training (Li et al., 2019; Reddi et al., 2019; Xu et al., 2019), thus improving model robustness and performance. In Figure 4, we illustrate the average gradient of item embeddings throughout the training process with and without our RepPad. We show the frequency of the gradient values as histograms. We can observe that the average gradient values are more centralized with the addition of our method. While the gradient is more likely to have large or small values during the original model training, RepPad mitigates this phenomenon, making the gradient more stable during the training process. Smooth updating of the gradient allows training to converge stably and improves the model’s recommendation performance
Figure 4. Average Gradient Histogram for Item Embeddings During Training. The ‘w/ RP’ represents adding our method to the corresponding model.
@@ 4.3. Summary and Insights
RepPad can make the gradient update of the model more stable without increasing the total number of training sequences. Unlike existing methods, our method augments short sequences to a length comparable to that of long sequences in a nearly lossless manner. Existing methods inevitably generate noise or compromise the integrity of the original data when perform augmentation. From another perspective, RepPad enables the model to focus fairly on short and long sequences, giving more attention to long-tail items and users. In this way, the model can better learn the representation of cold-start users and items. We also speculate that RepPad acts as some kind of regularization, which mitigates the negative effects of the uneven distribution of original dataset.
5. Related Works
5.1. Sequential Recommendation
Sequential recommendation (SR) is an emerging topic in the field of recommender systems (Qu et al., 2022). An early solution is to treat the item sequence as a Markov Chain (Rendle et al., 2010; Shani et al., 2005; He and McAuley, 2016), where the next item to predict is closely related to the latest few interactions. The limitation lies in the inability to learn the dependency in a relatively large time step. Hence, a better solution based on recurrent neural networks (RNNs) (Schuster and Paliwal, 1997) emerges since it can capture long- and short-term preferences. A pioneering work GRU4Rec (Hidasi et al., 2015a) introduced RNN into the field of recommender systems. It trained a Gated Recurrent Unit (GRU) architecture to model the evolution of user interests. Following this idea, CmnRec (Qu et al., 2022) proposed to divide proximal information units into chunks, whereby the number of memory operations can be significantly reduced. More recently, transformer-based approaches have attracted much attention because of their convenience for parallelization and carefully designed self-attention architecture (Vaswani et al., 2017; Devlin et al., 2018). SASRec (Kang and McAuley, 2018) employed unidirectional Transformers to fulfill the next-item prediction task in the sequential recommendation. Some works further optimized the Transformer’s architecture to make it more suitable for recommendation scenarios, such as reducing computational complexity (Fan et al., 2021) and stochastic self-attention (Fan et al., 2022). Furthermore, FMLP-Rec (Zhou et al., 2022) proposes an all-MLP architecture with learnable filters to enhance recommendation performance.
5.2. Data Augmentation for SR
Data augmentation has been established as a standard technique to train more robust models in many domains. In recommender systems, Slide Windows is an early work (Tan et al., 2016b), which modified the original click sequence by splitting it into many sub-sequences or randomly discarding a certain number of items. Inspired by counterfactual thinking, CASR (Wang et al., 2021) revised the sequence of user behaviors by substituting some previously purchased items with other unknown items. Some researchers proposed to leverage bi-directional transformers (Liu et al., 2021b; Jiang et al., 2021) and diffusion models (Liu et al., 2023) for sequence data augmentation. With the rise of self-supervised techniques (e.g., contrastive learning), many researchers used it as an auxiliary task to further improve model performance (Qin et al., 2023; Xie et al., 2022). For example, CL4SRec (Xie et al., 2020) developed three data augmentation operators for pre-training recommendation tasks equipped with a contrastive learning framework. Based on CL4SRec, CoSeRec (Liu et al., 2021a) integrated item similarity information for data augmentation with a contrastive learning objective, aiming to maximize the agreement of augmented sequences. Unlike existing methods, we are the first to explore the use of idle padding space for data augmentation. In terms of simplicity and applicability, our method does not contain any parameters or training processes and can be plugged into most existing sequential models.
@@ 6. Conclusion
In this work, we start from the commonly used padding strategy, aiming to utilize the idle input space to further improve the model performance and training efficiency. Based on this idea, we propose RepPad, which leverages the original sequence as padding content instead of the special value 0. The RepPad can be considered as a sequence-level data augmentation method. Compared with existing methods, our method does not contain any parameters or training processes, and it is a data augmentation plugin with high applicability. Extensive experiments demonstrate the superiority of RepPad. It improves the performance of various sequential models and outperforms existing data enhancement methods. We also provide multiple perspectives and insights into what makes RepPad effective. For future work, we hope to provide theoretical underpinnings for RepPad or explore more appropriate padding strategies. We are also interested in combining RepPad with other data augmentation methods.
Acknowledgements.
This work is partially supported by the National Natural Science Foundation of China under Grant No. 62032013, the Science and technology projects in Liaoning Province (No. 2023JH3/10200005), and the Fundamental Research Funds for the Central Universities under Grant No. N2317002.

### v3

@@ 4.5. Ablation Study
We conduct an ablation study to explore the different variants of RepPad and RepPad+. Based on the padding times $m$ in Table 1 and whether or not to add the delimiter 0, we compare the recommendation performance of the following variants: 1) Fix: Performing repeated padding with a fixed number of times, we report the best results for times in range $\{1,2,3\}$. 2) Max: Performing repeated padding until the remaining space is insufficient (i.e. maximum padding times). 3) (0, Max): When performing repeated padding each time, randomly select between 0 and the maximum number of times. 4) (1, Max): When performing repeated padding each time, randomly select between 1 and the maximum number of times. Figure 3 and 4 presents the variants and results. We report the performance of each of the four variants with and without the addition of the delimiter 0.
Figure 3. The result of the ablation study for RepPad.
Figure 4. The result of the ablation study for RepPad+.
@@
Compared to the original model, RepPad+ delivers more overall performance gains than RepPad. Looking at the different variants, Fix and Max bring limited performance improvement compared to (0, Max) and (1, Max). The performance of (0, Max) and (1, Max) can come with considerable improvements. Further, (1, Max) performs slightly better than (0, Max) on the whole, indicating that performing repeat padding for short sequences at least once is essential. For Fix, our methods are performed with a fixed number of times, so it does not fully utilize the idle input space. For Max, we observe that it not only does not bring improvement in some cases but also leads to a decrease in model performance. If every sequence performs repeated padding to the maximum number of times, it may make the representation learned by the model unbalanced. Because the repeated padding time tends to be large for short sequences, multiple iterations of short behavioral sequences may interfere with the model’s ability for long-sequence and long-term preferences modeling, which in turn impairs recommendation performance. For (0, Max) and (1, Max), they avoid the imbalance problem caused by RP-Max while using the idle input space wisely.
@@
For whether or not to add the delimiter 0, we observe that this factor is related to the type of recommendation model in Figure 3. For GRU4Rec, there are more examples of better performance when no delimiter is added. In contrast, for SASRec, there are more examples that models perform better when the delimiter is added. In other words, RNN-based models are less susceptible to the interference of ”predicting the tail with the head”. The self-attention module of Transformer-based models pays attention to the effect of each historical interaction on the following action, so the negative effect of ”predicting the tail with the head” can be weakened by adding the delimiter 0. In Figure 4, these trends are just the opposite, i.e., GRU4Rec performs better with the addition of delimiter 0 and SASRec performs better without the delimiter 0. In addition, we observe that the need to add the delimiter 0 is also relevant to the dataset. For example, for variant (1, Max), the trend of performance advantages and disadvantages with and without 0 is very similar when using different backbone networks under the same dataset. As the only adjustable option in our method, we recommend users try adding 0 and not adding it respectively in real applications.
Table 10. The training speed (seconds) of RepPad+ and different baseline methods. We give the increased time compared to the original model. The best performance is bolded, and the second best is underlined.
Dataset GRU4Rec SASRec
Method Speed (s/epoch) Total Time (s) H@10 Method Speed (s/epoch) Total Time (s) H@10
Beauty Original 0.960 153 0.0327 Original 2.389 395 0.0637
w/ CMRSI 2.547 (+1.587) 507 (+354) 0.0463 w/ CMRSI 4.561 (+2.172) 1416 (+1021) 0.0603
w/ CL4SRec 1.748 (+0.778) 329 (+176) 0.0440 w/ CL4SRec 3.954 (+1.565) 677 (+282) 0.0686
w/ Fix 1.103 (+0.143) 203 (+60) 0.0447 w/ Fix 2.574 (+0.185) 522 (+127) 0.0695
w/ (1,Max) 1.392 (+0.432) 168 (+15) 0.0505 w/ (1,Max) 2.806 (0.0417) 503 (+108) 0.0762
Sports Original 1.422 210 0.0143 Original 3.133 642 0.0299
w/ CMRSI 3.853 (+2.431) 1081 (+871) 0.0176 w/ CMRSI 6.400 (+3.267) 1066 (+424) 0.0318
w/ CL4SRec 2.620 (+1.198) 561 (+351) 0.0271 w/ CL4SRec 5.479 (+2.346) 928 (+286) 0.0339
w/ Fix 1.585 (+0.163) 248 (+38) 0.0204 w/ Fix 3.237 (+0.104) 746 (+104) 0.0376
w/ (1,Max) 2.087 (+0.655) 261 (+51) 0.0232 w/ (1,Max) 3.782 (+0.649) 861 (+219) 0.0434
Home Original 2.531 391 0.0057 Original 4.840 958 0.0155
w/ CMRSI 16.013 (+13.482) 1560 (+1169) 0.0094 w/ CMRSI 13.464 (+8.624) 2837 (+1879) 0.0173
w/ CL4SRec 5.101 (+2.570) 939 (+548) 0.0156 w/ CL4SRec 9.196 (+4.356) 1938 (+980) 0.0220
w/ Fix 2.786 (+0.255) 400 (+9) 0.0086 w/ Fix 5.156 (+0.316) 1077 (+119) 0.0189
w/ (1,Max) 3.745 (+1.214) 421 (+30) 0.0121 w/ (1,Max) 5.987 (+1.147) 1147 (+189) 0.0211
@@
5. Why Our Method Works
The experimental results in Section 4 show that our approach achieves improvements on various types of sequential models, including CNNs, RNNs, MLPs, and Transformers. It is difficult to get a generalized interpretation if we analyze from only a single model type. Therefore, we seek more generalized explanations regarding the common points in all the sequential models.
5.1. Training Efficiency
In Figure 5, we illustrate the loss convergence of different baseline methods and our RepPad+ during training. We also compare the epoch time consumption and total training time of different methods in Table 10. For a fair and intuitive comparison, we only record the time spent on the model training step. The time spent on the validation and eventual testing steps required for the early-stopping strategy is not included.
From Figure 5, we can observe that our methods’ losses converge slightly more stable and faster than the original model. For CMRSI and CL4SRec, the heuristic data augmentation introduces more training samples along with more noise, resulting in more epochs for the model to converge. The contrastive learning task of CL4SRec also imposes a convergence burden. Compared to the original model, Fix shows a slight improvement in convergence speed, while (1,Max) shows a significant improvement in all cases. From Table 10, we can observe that the two variants of RepPad+ only add a small amount of epoch time and total training time compared to the original model. The CMSRI and CL4SRec add extra training samples and auxiliary tasks, substantially increasing computational time. On the Home dataset, which has a large number of interactions, the training time increases by up to a factor of 3. While RepPad+ usually only brings 0.2 to 0.3 times extra training time. Comparing Fix and (1, Max), the former can only perform repeated padding a fixed and smaller number of times. The latter randomly chooses between one time and the maximum number of times, which slightly increases the time-consuming but further improves the loss convergence and model performance. In most cases, (1,Max) can achieve a win-win situation regarding training speed and model performance.
Figure 6. Average gradient value histogram for all item embeddings during training.
Table 11. Performance of three special variants of RepPad+.
Dataset Method GRU4Rec SASRec
H@10 N@10 H@20 N@20 H@10 N@10 H@20 N@20
Beauty Original 0.0327 0.0162 0.0551 0.0219 0.0637 0.0331 0.0942 0.0408
w/ Ours-$\alpha$ 0.0312 0.0157 0.0530 0.0197 0.0618 0.0322 0.0903 0.0382
w/ Ours-$\beta$ 0.0284 0.0144 0.0464 0.0188 0.0622 0.0315 0.0912 0.0394
w/ Ours-$\gamma$ 0.0315 0.0158 0.0495 0.0203 0.0533 0.0304 0.0776 0.0365
Sports Original 0.0143 0.0065 0.0249 0.0091 0.0299 0.0160 0.0447 0.0197
w/ Ours-$\alpha$ 0.0127 0.0060 0.0216 0.0081 0.0284 0.0156 0.0413 0.0187
w/ Ours-$\beta$ 0.0132 0.0061 0.0225 0.0084 0.0289 0.0152 0.0426 0.0184
w/ Ours-$\gamma$ 0.0079 0.0038 0.0131 0.0052 0.0198 0.0098 0.0324 0.0130
@@ 5.2. Gradient Stability
Research has shown that more stable or smooth gradients are more helpful for model training (Li et al., 2019; Reddi et al., 2019; Xu et al., 2019), thus improving model robustness and performance. In Figure 6, we illustrate the average gradient of item embeddings throughout the training process with and without our RepPad+. We show the frequency of the gradient values as histograms. We can observe that the average gradient values are more centralized with the addition of our method. While the gradient is more likely to have extremely large or small values during the original model training, RepPad+ mitigates this phenomenon, making the gradient more stable during the training process. The RepPad+ increases the number of training samples input into the model, especially for short interaction sequences. The model can derive more stable gradient estimates from more samples, thus accurately modeling user preferences.
5.3. Other forms of RepPad+
In order to verify whether the effectiveness of RepPad+ is due to increasing the number of iterations in which the data is involved, we implemented RepPad+ in a different form. First, we completed a model training session using (1,Max), during which we recorded the number of times that each sequence was subjected to the execution of RepPad+ and saved it. Afterward, we retrain a model using traditional padding. However, the number of times each sequence is fed into the model at each epoch is determined by the number of times previously recorded. This variant is denoted by Ours-$\alpha$. We also use RepPad+ in the inference phase of the model instead of the training phase. This variant is denoted by Ours-$\beta$. In addition, we replace the padding content of repeated padding with other user sequences. This variant is denoted by Ours-$\gamma$.
The results are illustrated in Table 11. We can observe that neither using RepPad+ in the inference phase nor changing the number of input data improves the performance of the original model and even results in performance degradation. For Ours-$\alpha$, it treats each repetition of padding as feeding the original sequence into the model once, which changes our RepPad+’s idea of data augmentation from horizontal to vertical but produces a performance degradation compared to the original model. This proves that the use of idle input space impacts the model’s training process, which can be seen as adding a regularization to the training process. With this regularization, the model can more equitably focus on short and long sequences in the dataset, providing more accurate recommendations for users with short sequences and reducing the negative impact of long-tail distributions. For Ours-$\beta$, the change in the inference phase will not bring any new knowledge to the model. Given that the training data is distributed with the test data, using RepPad+ at this point will affect the data distribution in the inference phase, which in turn will jeopardize the model performance. From the performance of Ours-$\gamma$, we can observe that randomly padding other sequences leads to a significant degradation of the recommendation performance. This randomized padding strategy interferes with the model’s learning of sequence patterns contained in the original data, and the mixing of different sequences may cause the model to learn the incorrect user preferences. In our methods, padding the original sequence maximizes the preservation of that user’s preference patterns and avoids the above situation.
Figure 7. An illustration of information leakage problems for our methods.
Table 12. The results of future informative leakage analysis. The backbone network is SASRec.
Dataset Method H@10 N@10 H@20 N@20
Beauty Original 0.0644 0.0335 0.0961 0.0415
w/ RepPad+ 0.0762 0.0424 0.1089 0.0507
w/ RepPad+ and Avoid Leakage 0.0699 0.0385 0.1009 0.0467
Sports Original 0.0299 0.0160 0.0447 0.0197
w/ RepPad+ 0.0369 0.0196 0.0553 0.0242
w/ RepPad+ and Avoid Leakage 0.0346 0.0180 0.0523 0.0225
5.4. The Leakage of Future Information
Most sequential recommendation models can be viewed as autoregressive models. They predict the next interaction based on the user’s historical sequence data. When training or testing autoregressive models, it is usually necessary to avoid leakage of future information, i.e., the model sees in advance what needs to be predicted when making a prediction. Usually, future information leakage during training can lead to a decrease in the performance and reliability of the model. However, as shown in Figure 7, we observe that after using our method, there is a leakage of future information, i.e., the model would use future interactions to predict past interactions during training. Also, adding the delimiter 0 does not prevent ”predicting the tail with the head”. It can be considered as attenuation of attention or influence because the items in the sequence are relatively farther apart after adding the delimiter 0.
Counterintuitively, the above information leakage does not degrade the model performance. It may bring more performance gains to the model than the negative impacts it causes. We mask the red line in Figure 7 during training, and the results are presented in Table 12. We can observe that the model performance decreases after blocking the effect of future information leakage, which proves to be beneficial for recommendation accuracy. Some researchers have also explored using future information to improve model performance (Yuan et al., 2020). In sequential recommendation, although user behavior is in the form of sequential data, sequential dependencies may not be strictly maintained. For example, after a user purchases a cell phone, she may purchase a protective case, headphones, or a screen protector, but there is no sequential dependency between these three items. In other words, the user will likely click on these three items in any order. Our method weakens the notion of strict sequence order in previous sequential recommendations.
@@ 5.5. Summary and Insights
Based on the analysis above, we believe that our method utilizes the idle input space to provide more samples for model training and improve the training efficiency. Our method can make the gradient update of the model more stable without increasing the total number of training sequences. Unlike existing methods, RepPad augments short sequences to a length comparable to that of long sequences in a nearly lossless manner. RepPad+ improves RepPad by augmenting medium-length sequences, further improving the efficiency of idle space utilization. Existing methods inevitably generate noise or compromise the integrity of the original data when performing augmentation. From another perspective, our method enables the model to focus fairly on short and long sequences, giving more attention to long-tail items and users. In this way, the model can better learn the representation of cold-start users and items and thus give more accurate recommendations. This is similar to L2Aug (Wang et al., 2022), which categorizes users into core and casual users. The core users tend to have more interaction data, while casual users tend to have short interaction sequences. It uses the historical data from the former to augment the interaction sequence of the latter. However, our method does not include any training process and parameters and does not increase the total amount of training data. Furthermore, utilizing future information is also an interesting finding, similar to the idea presented in GRec (Yuan et al., 2020) that future information provides valuable signals about user preferences and can be used to improve the quality of recommendations. How to utilize them wisely is also a valuable research direction. We also speculate that RepPad and RepPad+ act as some kind of regularization during model training, which mitigates the negative effects of the uneven distribution of the original dataset.
@@ 6. Conclusion
In this work, we argue that the existing sequential recommendation’s padding strategy results in a large amount of wasted input space, which can be further utilized to improve the model performance and training efficiency. Based on this idea, we propose an extremely simple yet effective data augmentation method, RepPad+, which leverages the original sequence as padding content instead of the special value 0. Our method contains no hyper-parameters or learnable parameters compared to existing data augmentation methods. It is a data augmentation plugin with high applicability and generalization. Extensive experiments on multiple datasets demonstrate the superiority of our method. It improves the recommendation performance of various sequential models and outperforms existing data augmentation methods. In addition to recommending performance leadership, our approach is highly competitive in terms of efficiency. We also provide multiple analyses and perspectives into what makes RepPad+ effective. For future work, we plan to explore more effective padding strategies for sequential recommendation. We are also interested in combining RepPad+ with other data augmentation methods.
Acknowledgments
This work is partially supported by the National Natural Science Foundation of China under Grant No. 62032013, No. 62102074, and the Science and technology projects in Liaoning Province (No. 2023JH3/10200005).